OpenStudy (calculusxy):

The perimeter of an isosceles right triangle is 24 cm. How long are its sides?

OpenStudy (calculusxy):

@jim_thompson5910

OpenStudy (calculusxy):

@Nnesha @skullpatrol @mathmale

OpenStudy (seratul):



OpenStudy (calculusxy):

I thought of setting up an equation: \(a + a + a\sqrt{2} = 24\)

OpenStudy (seratul):

What do you know about the hypotenuse?

OpenStudy (calculusxy):

This would be a 45-45-90 triangle I suppose

OpenStudy (calculusxy):

The length would be \(a\sqrt{2}\)?

OpenStudy (seratul):

Simpler than that. Just in general, what do you know about it?

OpenStudy (calculusxy):

It's the longest side of the triangle

OpenStudy (seratul):

Correct. By the way, is this multiple choice? I am pretty sure there is more than one answer.

OpenStudy (calculusxy):

Nope.

OpenStudy (calculusxy):

@jim_thompson5910 That's where I am having the problem.

OpenStudy (seratul):

Welp, you got this jim :D

OpenStudy (calculusxy):

I only get up \(2a + a\sqrt{2} = 24\). I don't know whether I should factor it to \(a(2 + \sqrt{2}) = 24\) and if I do then how would I solve for \(a\)?

jimthompson5910 (jim_thompson5910):

why not factor out the GCF 'a' to go from \[\Large a+a+a\sqrt{2} = 24\] to \[\Large a(1+1+\sqrt{2}) = 24\]

jimthompson5910 (jim_thompson5910):

yes then you'd have \[\Large a(2+\sqrt{2}) = 24\]

OpenStudy (calculusxy):

Would I divide both sides by \((2 + \sqrt{2})\)?

jimthompson5910 (jim_thompson5910):

essentially the left side is saying `some number 'a' times (2+sqrt(2))`

jimthompson5910 (jim_thompson5910):

yes you would divide both sides by (2+sqrt(2))

OpenStudy (calculusxy):

So now I would have \(a = \frac{24}{(2 + \sqrt{2})}\)

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (calculusxy):

That's the complete answer? Like no more simplifying?

jimthompson5910 (jim_thompson5910):

you can rationalize the denominator if you want

jimthompson5910 (jim_thompson5910):

or convert to decimal form to get an approximate answer

OpenStudy (calculusxy):